express cos^(3)tetasin^(4)teta in terms of multiple angle
1
Expert's answer
2020-05-08T19:00:49-0400

in terms of complex numbers

let "z=\cos\theta+i\sin\theta"

then

"2\cos\theta=z+\frac{1}{z}\\2i\sin\theta=z-\frac{1}{z}"

then

"(2\cos\theta)^{3}(2i\sin\theta)^{4}=(z+\frac{1}{z})^{3}(z-\frac{1}{z})^{4}\\128\cos^{3}\theta\sin^{4}\theta=(z-\frac{1}{z})(z^{2}-\frac{1}{z^{2}})^{3}\\128\cos^{3}\theta\sin^{4}\theta=(z-\frac{1}{z})(z^{6}-3z^{2}+\frac{3}{z^{2}}-\frac{1}{z^{6}})\\128\cos^{3}\theta\sin^{4}\theta=(z^{7}+\frac{1}{z^{7}})-(z^{5}+\frac{1}{z^{5}})-3(z^{3}+\frac{1}{z^{3}})+3(z+\frac{1}{z})"

substituting back:

"128\cos^{3}\theta\sin^{4}\theta=2\cos7\theta-2\cos5\theta-3(2\cos3\theta)+3(2\cos\theta)"


dividing through by 128


"\\\cos^{3}\theta\sin^{4}\theta=\frac{1}{64}[\cos7\theta-\cos5\theta-3\cos3\theta+3\cos\theta]"


"\cos^{3}\theta\sin^{4}\theta=\frac{1}{64}[3\cos\theta-3\cos3\theta-\cos5\theta+\cos7\theta]"